IT the polar ofthe point (x, y) are Cr, e). (e) (ax, -ar the Cartesian

hello, please helo with thesebquestions. thanks
Answer #1

a)

r = sqrt ( 1^2+ 1^2) = 1.414

theta = 180 - arctan ( - 1/1) = 135

(r, theta) = ( 1.414, 135)

========

b)

r = sqrt [ (-2)^2+ (-2)^2 ] = 2.828

theta = 180 + arctan ( -2 / (-2)) = 225

(r, theta) = ( 2.828, 225)

========

c)

r = sqrt ( 3^2 + 3^2) = 4.24

theta = 360 - arctan ( 3/3) = 315

=======

do comment in case any doubt, will reply for sure.. that